### v2

# SimuRA: A World-Model-Driven Simulative Reasoning Architecture for General Goal-Oriented Agents

Abstract
AI agents built on foundation models hold enormous promise. Current practice, however, focuses on a one-task-one-agent approach, which not only falls short of scalability and generality, but also faces practical limitations from black-box autoregressive reasoning, where decisions unfold token by token without explicit simulation or counterfactual evaluation of outcomes. Humans, on the other hand, reason and plan by mentally simulating the consequences of actions within an internal model of the world – a capability that supports flexible, goal-directed behavior across diverse contexts. Moving towards a more general and powerful AI agent, we introduce SimuRA, a goal-oriented architecture for generalized agentic reasoning. Based on a principled formulation of an optimal agent in any general environment, SimuRA addresses the limitations of black-box autoregressive reasoning by incorporating the world model for planning via simulation. Our prototype world model is implemented using LLMs as a substrate, leveraging the natural language as a discrete, hierarchical representation grounded in concepts for planning, while remaining model-agnostic. On complex web-browsing tasks such as flight search, SimuRA improves the success rate from 0% to 32.2% compared to a representative open-web agent baseline. Across tasks, world-model-based planning achieves up to 124% higher task completion rates than a matched black-box autoregressive baseline, demonstrating the advantages of simulative reasoning. [^1] All experiments were conducted in 2024 with then-available models and tooling. We release ReasonerAgent-Web, a web-browsing agent built on SimuRA, as an open-source research demo.
1 Introduction
AI agents powered by foundation models such as Large Language Models (LLM) or Vision Language Models (VLM) have demonstrated tremendous potential for handling tasks that require flexible decision making. For instance, there have been releases of agents specialized in web and computer automation openai_cua_2025 ; deepmind2024mariner ; zhou2023webarena ; xie2024osworldbenchmarkingmultimodalagents , internet research openai_deepresearch_2025 ; anthropic2024claude35sonnet ; google_gemini_deep_research_2024 , social simulation park2023generativeagentsinteractivesimulacra , software development anysphere2025cursor ; wang2025openhandsopenplatformai , scientific research gottweis2025towards ; yamada2025aiscientistv2workshoplevelautomated , and so on. Despite the promise, these agent architectures remain tailored to specific tasks, limiting their scalability and transferability. Recent reasoning LLMs optimized via end-to-end reinforcement learning (RL) openai_learning_to_reason ; guo2025deepseek have begun to show traces of emergent planning and improved long-horizon control, suggesting that black-box autoregressive reasoning can partially internalize aspects of deliberative reasoning.  [^2] In this work, we use “autoregressive reasoning” to denote the procedural mode of decision-making that iteratively predicts the next abstract internal state $z_{t}$ (e.g., text or latent tokens) based on distribution $p(z_{t}\mid z_{<t},x)$ given input $x$, without explicit modeling or simulation of future outcomes – distinct from autoregressive modeling as a statistical or architectural property. However, their reasoning still proceeds primarily through token-by-token generation, lacking an explicit mechanism for simulating and evaluating counterfactual futures. As a result, even strong models can remain vulnerable to locally myopic decisions or error accumulation over extended trajectories andreas2022languagemodelsagentmodels ; yao2023react . Humans, in contrast, are capable of reasoning and planning to achieve goals in diverse environments. Using a single cognitive system, human beings adapt to different task and environments not only by linear reasoning, but also by imagining potential outcomes, simulating possibilities using a mental world model, and planning accordingly Ball_2020 ; lecun_path_2022 .
Figure 1: Demo of tasks performed using a web-browsing agent built on SimuRA with simulative planning using a world model.
Moving towards more powerful and generally applicable AI agents, we introduce SimuRA (Simulative Reasoning Architecture), a goal-oriented architecture for generalized agentic reasoning. SimuRA mitigates the limitations of linear autoregressive reasoning by introducing world model as the engine for planning via simulation. Specifically, a policy module first proposes a few potential actions, aimed at achieving specific goals based on agent identity and environment. The world model then simulates the outcomes of those proposed actions. Finally, a critic module evaluates these outcomes against the initial goals in order to select the best action from the candidates. Because simulating the full details of the world is infeasible and unnecessary for planning, we extract only the relevant information using natural language as a compact, semantically structured representation space for reasoning. To ensure robustness to observation noise and distracting execution details, we further propose a hierarchical architecture that isolates perception, simulative planning, and action selection, enhancing adaptability and consistency across diverse tasks.
We conducted experiments by implementing these modular components within a LLM-based pipeline, applied to a range of web browsing tasks such as complex website navigation, multi-hop, multi-website QA, and general web automation. In particular, to evaluate the ability of agents in robustly interacting with complex websites, we develop FlightQA (Section 4.1), a dataset of flight search questions with varying number of constraints and types of constraints (e.g., dates, transfer, pricing) that are time-insensitive and automatically verifiable. Results show that SimuRA performs substantially better compared to baselines, increasing the success rate on FlightQA from 0% to 32.2% compared to a popular open-web agent. In particular, world-model-based planning outperforms a matched autoregressive reasoning baseline by up to 124% in task-completion rate, showing the advantage of simulative reasoning. All experiments were conducted in 2024 with the then-available models and browser tooling. Figure 1 shows examples of the agent performing multi-website, long-range task such as flight searching, online shopping, and news research.
For evaluation and demonstration purposes, we implemented SimuRA as an open-source library available via LLM-Reasoners hao2024llm . [^3] https://github.com/maitrix-org/llm-reasoners/tree/main/examples/ReasonerAgent-Web The resulting web agent, ReasonerAgent-Web reasoneragent2025 , is available as an open-source research demo. [^4] https://easyweb.maitrix.org/ We are actively developing next versions of the system to address broader challenges and to further demonstrate its generality across a wider range of tasks and environments.
2 Related Work
LLM-Based Agents
LLM-based agents have rapidly evolved into versatile systems capable of autonomous behavior across a range of environments. One major approach to build such systems focuses on data collection in the targeted environment followed by model training. Notable examples include AutoWebGLM lai2024autowebglmlargelanguagemodelbased , AgentQ putta2024agentqadvancedreasoning , UI-TARS qin2025uitarspioneeringautomatedgui , etc. Prompt-based workflows, on the other hand, have also shown strong potential when equipped with carefully designed modules, as demonstrated by recent work such as AWM wang2024agentworkflowmemory , VOYAGER wang2023voyageropenendedembodiedagent , and so on. SimuRA is built on prompt-based workflows but can leverage observation data for targeted improvement of its world model chae2024web , leading to reduced reliance on human demonstration and strong generalizability to new tasks lecun2022pathtowards , which is an exciting next step.
World-Model-Based Agents
Model-based planning for agents have long been frequently discussed and studied. Early work demonstrated the success of this approach by testing in classic games like go, chess, shogi and Atari. oh2015actionconditionalvideopredictionusing ; Schrittwieser_2020 . Later on, world model was used for policy optimization and experimented on control tasks. janner2021trustmodelmodelbasedpolicy ; hansen2022temporaldifferencelearningmodel In recent years, with the boost in foundation model’s capabilities, world-model-based planning was applied to more complex problems like math reasoning hao2023reasoninglanguagemodelplanning , playing Minecraft hafner2024masteringdiversedomainsworld , and web browsing gu2025llmsecretlyworldmodel . However, these world models typically represent and/or predict the world states using holistic continuous embeddings, which suffer from noise and high variability which detracts from robust and stable decision-making barrett2017emotions . SimuRA instead adopts natural language as a discrete, concept-based latent space for consistent representation and prediction, which shows more general applicability across tasks in practice.
Web Browsing Agents
Web browsing and navigation were chosen to evaluate SimuRA due to their realism and the complex decision-making they demand across diverse, dynamic interfaces. Recent years have seen the emergence of several prominent web-browsing agents, from proprietary ones such as OpenAI’s Operator openai_cua_2025 , Anthropic’s Computer Use anthropic2024claude35sonnet , and Google-DeepMind’s Project Mariner deepmind2024mariner , to open-source ones including OpenHand’s BrowsingAgent openhands , WebVoyager he2024webvoyagerbuildingendtoendweb , CogAgent hong2024cogagentvisuallanguagemodel and WebAgent gur2024realworldwebagentplanninglong . These agents are typically built on simple ReAct-based yao2023react autoregressive reasoning which can have difficulty recovering from previous mistakes; their often specialized design also preclude these approaches from generalizing to other task domains like social interactions and the physical world. Numerous benchmarks have been introduced to evaluate these web agents, including WebArena zhou2023webarena , WebVoyager he2024webvoyagerbuildingendtoendweb MiniWoB++ liu2018reinforcement , Mind2Web deng2023mind2webgeneralistagentweb , and WebShop yao2023webshopscalablerealworldweb . Despite wide adoption, these benchmarks are usually either built in simulated and simplified environments, based on outdated questions, or lacks convincing method of measuring task completion, which detract from the goal of evaluating practically useful web agents. To address these challenges, we build FlightQA, an new dataset for evaluating agent ability in real-time complex website navigation. More details are included in Section 4.1.
Generalist Agents
There have been various attempts of building generalist agents recently. One class of methods focuses on building multi-agent workflows where different tasks or subtasks are delegated to different specialist agents hu2025owloptimizedworkforcelearning ; jiang2025longtermmemoryfoundation ; zhang2025agentorchestrahierarchicalmultiagentframework ; fourney2024magenticonegeneralistmultiagentsolving . While versatile and extensible, improvements of the system often requires implementation of new specialist agents. We aim to build a holistic architecture capable of leveraging shared experience across different domains and modalities, a core benefit enabled by world model learning. Alternatively, there is another major line of work focus on composing executable scripts to flexibly control predefined tools, thus enabling generalist behavior wang2024executablecodeactionselicit ; openhands ; smolagents ; qiu2025alitageneralistagentenabling . Our goal is to further enhance the flexibility by allowing the agent to make high-level plans with natural language, think ahead, update its plans, and correct potential mistakes during execution.
3 SimuRA: Generalized Architecture for Optimal Goal-Oriented Agent
3.1 Formulation of Agent-Environment Model
We first present our formulation of an optimal goal-oriented agent following the agent-environment model presented in xing2025critiques : We consider an agent $\pi$ with identity $i$ (e.g., name, description) and goal $g$ acting in environment $\mu$ (e.g., web browser, physical world, the entire universe) with action space $\mathcal{A}$ and state space $\mathcal{S}$ (Figure 2). Formally, at each time step $t$, the agent $\pi$ takes the current state $s_{t}\in\mathcal{S}$ and outputs the next action $a_{t}\in\mathcal{A}$ following a policy distribution $p_{\pi}(a_{t}\mid s_{t})$, while the environment $\mu$ takes the current state $s_{t}$ and action $a_{t}$, and outputs the next state $s_{t+1}\in\mathcal{S}$ based on the distribution $p_{\mu}(s_{t+1}|s_{t},a_{t})$. We can thus denote the distribution of the interaction trajectory up to timestep $T$, or $(a_{t},s_{t+1},\dots,a_{T-1},s_{T})$ given the current state $s_{t}$, as below:
$p^{\pi}_{\mu}(a_{t},s_{t+1},\dots,a_{T-1},s_{T}\mid s_{t})=\prod_{k=t}^{T-1}{\underbrace{\textstyle p_{\pi}(a_{k}\mid s_{k})}_{\text{ agent }}}\ {\underbrace{\textstyle p_{\mu}(s_{k+1}\mid s_{k},a_{k})}_{\text{ environment }}}$ (1)
In each state $s_{t}$, the agent also receives a reward $r(g,s_{t})$ based on its goal $g$. We evaluate the agent by its discounted cumulative reward, denoted as $\sum_{k=t}^{\infty}\gamma_{k}r(g,s_{k})$ (with the discount parameter $\gamma_{t}$ decaying to zero with time, i.e., $\lim_{t\to\infty}\gamma_{t}=0$). Note that this reward function can be dense (e.g., gaming scores), but perhaps frequently sparse (e.g., curing a disease). The agent’s long-term success can thus be measured by its expected future discounted reward, also known as value function sutton1998reinforcement , which satisfies the following recurrence:
$V_{\pi,\mu}^{g}(s_{t})$ $\vcentcolon=\mathbb{E}_{\pi,\mu}\left[\sum_{k=t}^{\infty}\gamma_{k}r(g,s_{k})\mathrel{\bigg|}s_{t}\right]$
$=\lim_{T\rightarrow\infty}\sum_{(a_{t},s_{t+1},\dots,s_{T})}\sum_{k=t}^{T}\gamma_{k}r(g,s_{k})\ p^{\pi}_{\mu}(a_{t},s_{t+1},\dots,s_{T}\mid s_{t})$
$=\sum_{(a_{t},s_{t+1},\dots,s_{T})}\Bigg({\underbrace{\sum_{k=t}^{T-1}\gamma_{k}r(g,s_{k})+\gamma_{T}V_{\pi,\mu}^{g}(s_{T})}_{\text{goal progress}}}\Bigg)\ {\underbrace{\textstyle p^{\pi}_{\mu}(a_{t},s_{t+1},\dots,s_{T}\mid s_{t})}_{\text{trajectory}}},$ (2)
Which indicates that the value function in state $s_{t}$ can be expressed in terms of the value function at possible future states $s_{T}$ weighted by their probabilities.
Figure 2: A possible definition of an optimal agent
3.2 Definition of Optimal Agent
Based on Equations 1 and 2, we can define the optimal agent $\pi^{*}_{\mu}$ in environment $\mu$ as one that maximizes the value function, written formally as below:
$\pi^{*}_{\mu}:=\operatorname*{arg\,max}_{\pi}V^{g}_{\pi,\mu}.$ (3)
Some simple derivation will show that the optimal agent in state $s_{t}$ will follow the following decision rule $\pi^{*}_{\mu}$ when planning for actions $a_{t:T-1}$:
$\pi^{*}_{\mu}(s_{t})={\underbrace{\operatorname*{arg\,max}_{a_{t:T-1}}}_{\text{possible actions}}}\ \sum_{s_{t+1:T}}\Bigg({\underbrace{\sum_{k=t}^{T-1}\gamma_{k}r(g,s_{k})+\gamma_{T}V_{\pi,\mu}^{g}(s_{T})}_{\text{goal progress}}}\Bigg)\prod_{i=t}^{T-1}\ {\underbrace{p_{\mu}(s_{i+1}\mid s_{i},a_{i})}_{\text{universe response}}}\$(4)
In practice, agents often samples promising action candidates using a policy function $\tilde{\pi}$ through the distribution $p_{\tilde{\pi}}(a_{t}\mid s_{t})$. Building the optimal agent thus requires capabilities for proposing possible actions ($\tilde{\pi}$), predicting their outcomes ($\mu$), and evaluating goal progress ($r,V$), respectively. Note that typical reactive agents that output the next action directly can be seen as taking the first sample from $\tilde{\pi}$ (similar to “System 1” in humans which makes fast, instinctive reactions kahneman2011thinking ), without simulating and evaluating the outcomes using $\mu$ and $V$ (similar to “System 2” responsible for deliberate decision-making), which provide opportunities for spotting and correcting errors from the deliberation process.
3.3 World Model for Generalized Simulative Reasoning
Note that the optimal decision-making process defined in Equation 4 requires the agent to have access to the ground-truth world state $s$ and the environment $\mu$ to experience and optimize over. However, these are often not available aside from simple scenarios like Go and Chess games (silver2016mastering, ; silver2017mastering, ) – imagine building a spacecraft to land on Mars, or simply a humanoid robot relying on noisy sensors in daily environments. World Model (WM) thus arises as a crucial component for predicting any environment’s response to a general agent. Specifically, a WM $f$ operates on an internal representation of the world state, denoted as a belief state $\hat{s}_{t}$, which is derived from sensory inputs $o_{t}$ via an Encoder $h$ (unlike the optimal agent described in §3.2 which has direct access to the true world state $s_{t}$). Given proposed action $a_{t}$, the WM predicts the next belief state $\hat{s}_{t+1}$ according to the distribution $p_{f}(\hat{s}_{t+1}\mid\hat{s}_{t},a_{t})$. The predicted belief state then allows the agent to propose the next action, continuing the cycle of prediction and action up to the desired time horizon $T$. Thus, a WM here essentially functions as a generative model of possible future world states, which enables simulative reasoning, or “thought experiments". Formally, for the optimal agent $\pi^{*}_{f}$ equipped with WM $f$ in belief state $\hat{s}_{t}$, we define the simulation-based decision rule in Equation 6 as follows:
$\pi^{*}_{f}(\hat{s}_{t})={\underbrace{\operatorname*{arg\,max}_{a_{t:T-1}}}_{\text{possible actions}}}\ \sum_{\hat{s}_{t+1:T}}\Bigg({\underbrace{\sum_{k=t}^{T-1}\gamma_{k}r(g,\hat{s}_{k})+\gamma_{T}V_{\pi,f}^{g}(\hat{s}_{T})}_{\text{goal progress}}}\Bigg)\prod_{i=t}^{T-1}\ {\underbrace{p_{f}(\hat{s}_{i+1}|\hat{s}_{i},a_{i})}_{{\scriptsize\shortstack{simulation with\\ world model}}}}$ (5)
A general-purpose WM $f$ here enables simulation of diverse possibilities across a wide range of domains, enabling agents to reason about outcomes without direct interaction with the environment.
3.4 Design of Simulative Reasoning Agent with Concept-based Latent States and Hierarchical Planning
In this subsection, we present our design of a generalizable simulative reasoning agent architecture. In particular, we provide detailed discussion on design decisions that enable robust and broad applicability across environments and tasks.
Discrete, Hierarchical State Representation via Natural Language
The dominant approach to encoding observation $o_{t}$ (e.g., webpages, video streams) has been to directly mapping all input tokens into continuous embeddings with fixed dimensionalities $\hat{s}^{z}_{t}$. While technically preserving all information, real-world sensory readings often suffer from inherent noise and high variability (e.g., ads on a webpage, varying weather and lighting conditions in video), which can make them brittle to reason over. Human cognition, on the other hand, has evolved to counter this variability by categorizing raw perception into discrete concepts barrett2017emotions , which are often encoded in language, symbols or structured thoughts. Indeed, natural language is inherently hierarchical, capable of encoding concepts from concrete ones (e.g., apple) to highly abstract ones (e.g., religion). Discrete representations are also complete in general xing2025critiques , which ensures no information is necessarily lost in the compression process. Implementing this form of perception, we propose to represent the world state $\hat{s}_{t}$ using a discrete natural language summary $\hat{s}^{c}_{t}$ generated by a pretrained encoder model $h$, formally expressed as below:
$p_{h}(\hat{s}_{t}\mid o_{t})=\prod_{i=1}^{N_{t}}p_{h}(\hat{s}_{t,i}\mid\hat{s}_{t,<i},o_{t}),$ (6)
Where each $\hat{s}_{t,i}$ is a natural language token. Likewise, we also denote the WM $f$ which predicts the next state $\hat{s}_{t+1}$ as a natural language sequence $\hat{s}^{c}_{t+1}$, formally as below:
$p_{f}(\hat{s}_{t+1}\mid\hat{s}_{t},a_{t})=\prod_{i=1}^{N_{t+1}}p_{h}(\hat{s}_{t+1,i}\mid\hat{s}_{t+1,<i},\hat{s}_{t},a_{t})$ (7)
Such a concept-based representation allows the other modules like policy to operate on a more structured latent space, which we find empirically to reduce hallucination and enable more robust planning, leading to better task performance in practice.
Figure 3: An agent in real world where groundtruth world state and universe are unavailable to experience or experiment, so world model is crucial for simulation. As discussed in §3.4, separation of simulated actions $a_{t}^{\prime}$ for planning and concrete actions $a_{t}$ for execution facilitates transfer and hierarchical planning, leading to more diverse and grounded actions which lead to better task success.
Hierarchical Planning via Simulated Actions
The customary approach to decision-making with world models has been to perform simulations or rollouts based on the specific action space $\mathcal{A}(\pi)$ afforded to the agent. While this approach indeed captures all the execution details, the specific idiosyncrasies of individual action spaces (e.g., parameter ordering, format, and scale) may hinder the transfer of knowledge across different action spaces, environments, and tasks, thereby limiting generalizable reasoning. Indeed, the real world may contain a richer range of intentions than what a particular action space offers (e.g., clicking on a flight may mean either exploring the pricing or committing to the option). Last but not least, the sequential roll-out over atomic actions can be inefficient and increase opportunities for error accumulation across multi-step, low-level predictions (e.g., swooshing of liquids with each muscle twitch), when higher-level dynamics over more abstract actions (e.g., spilling water due to tilting the glass) remain stable and predictable. To close this gap, we adopt a hierarchical architecture which separates high-level, flexible planning from low-level, rigorous execution sutton1991dyna . As illustrated in Figure 3, the agent’s policy $p_{\tilde{\pi}}(a^{\prime}_{t}\mid\hat{s}_{t})$ and world model $p_{f}(\hat{s}_{t+1}\mid\hat{s}_{t},a^{\prime}_{t})$ operate over simulated actions $a^{\prime}_{t}$ from a separate action space $\mathcal{A}^{\prime}$, while another actor $p_{\alpha}(a_{t}\mid a^{\prime}_{t},\hat{s}_{t})$ is responsible for selecting the concrete action $a_{t}\in\mathcal{A}$ conditioned on the selected simulated action $a_{t}^{\prime}$. This divide-and-conquer approach allows for more generalized reasoning disentangled from the exact details of the concrete action space and enables representation of a richer set of intentions. Furthermore, each simulated action $a_{t}^{\prime}$ may represent multiple execution steps in the environment (e.g., “explore the website” vs “click on the link”), which shortens the number of rollout steps for higher efficiency and fewer chances for error accumulation. In practice, we represent simulated actions $a_{t}^{\prime}$ using natural language due to its generality and expressivity, and find it results in more diverse and grounded action proposals, leading to better task success.
Figure 4: Illustration of the design of simulative reasoning agent with concept-based latent states and hierarchical planning.
Having discussed our major designs, we proceed to describe the full decision process of the SimuRA architecture: As illustrated in Figure 4, given observation $o_{t}$ (e.g., webpage screenshots and/or accessibility tree), SimuRA first infers the world state $\hat{s}_{t}$ using the encoder $h$, and then selects the best simulated action $a_{t}^{\prime*}$ through the planner. Inside the planner, the architecture performs simulative reasoning by proposing actions $a_{t}^{\prime}$ using policy $\tilde{\pi}$ and predicting the next state $\hat{s}_{t+1}$ using the world model $f$, and evaluating goal progress $\sum_{k=t}^{T^{\prime}-1}\gamma_{k}r(g,\hat{s}_{k})+\gamma_{T^{\prime}}V_{\pi,f}^{g}(\hat{s}_{T^{\prime}})$ using critic $v$ upon reaching state $\hat{s}_{T^{\prime}}$ at the planning horizon $T^{\prime}$. This can repeat multiple times until the planner selects the action sequence $a^{\prime*}_{t:T^{\prime}-1}$ with the highest expected success and passes the first step $a_{t}^{*}$ to actor $\alpha$ which finally outputs the concrete action $a_{t}$. Formally, SimuRA can be seen as solving the following multi-level optimization problem:
$\hat{s}_{t}$ $=\operatorname*{arg\,max}_{\hat{s}}\ {\underbrace{p_{h}(\hat{s}\mid o_{t})}_{\text{encoder}}}$ (Perception)
$a_{t:T^{\prime}-1}^{\prime*}$ $={\underbrace{\operatorname*{arg\,max}_{a_{t:T-1}^{\prime}}}_{\scriptsize\shortstack{sampled from \\ policy $\tilde{\pi}$}}}\sum_{\hat{s}_{t+1:T^{\prime}}}\ {\underbrace{v(\hat{s}_{T^{\prime}})}_{\text{critic}}}\prod_{k=t}^{T^{\prime}-1}{\underbrace{p_{f}(\hat{s}_{k+1}\mid\hat{s}_{k},a_{k}^{\prime})}_{{\scriptsize\shortstack{world model}}}}$ (Planning) (8)
$a_{t}$ $=\operatorname*{arg\,max}_{a}\ {\underbrace{p_{\alpha}(a\mid\hat{s}_{t},a_{t}^{\prime*})}_{\begin{array}[]{c}\scriptsize\text{actor}\text{}\end{array}}}$ (Acting) (10)
In §4 below, we present an instance of SimuRA with each of these components implemented using pretrained LLMs. While these LLMs alone are often insufficient for many complex agentic tasks, SimuRA’s divide-and-conquer approach combines existing LLM strengths like instruction-following, summarization, reflection, and tool use to allow agentic behavior to emerge. Benefiting from massive web-scale pretraining on next-token prediction $p(x_{t}\mid x_{<t})$, which is formally akin to world modeling, LLMs possess significant potential to serve as world models with natural-language state and simulated action spaces hao2023reasoninglanguagemodelplanning ; hu2023language . We approximately infer the world state $\hat{s}_{t}$ and action $a_{t}$ by sampling from the LLM-based encoder and actor distributions $p_{h}$ and $p_{\alpha}$, respectively. For planning, we optimize over the sampled actions $a_{t:T^{\prime}-1}^{\prime}$ using readily available tree search algorithms like Depth-First Search (DFS) and Monte-Carlo Tree Search (MCTS).
4 Experiments
Our proposed SimuRA architecture is generally applicable to various environments and tasks. As our first step, we evaluate our implementation on web browsing as an example due to both its practical value and its technical challenge. Web browser is an indispensable portal for individuals to perform many tasks in real life (e.g., gather information, book travels, and submit applications). Whereas many existing products do access the internet (chatgpt_search, ; perplexity, , etc.), they typically use specialized tools (e.g., search engines and data APIs) to capture a subset of web browser capabilities (i.e., reading) while falling short of the full functionality (e.g., access content not exposed to search engines or predefined APIs like flight and hotel databases). We argue that an agent that takes advantage of the full browser will push the envelope in AI’s abilities to serve human needs.
Figure 5: LLM-based implementation of SimuRA for web-related tasks (e.g. multi-website QA, flight search, etc). Planner is where we implement our proposed world-model-based planning. We also implement a baseline that simply samples the plan from a language model (i.e., autoregressive planning).
Despite the richness and flexibility, the web browser is a highly challenging environment for agentic reasoning due to its immense complexity, long-horizon nature, partial observability, and multimodality (zhou2023webarena, ; gu2024your, ). We evaluate our architecture in 3 types of web browsing tasks: 1) complex website navigation, 2) multi-hop, multi-website QA, and 3) general web automation. For the baselines, we compare against:
  1. 1.
    BrowsingAgent from OpenHands (openhands, ), a representative open-web agent which generates chain-of-thought before selecting an action.
  2. 2.
    SimuRA (our architecture) with autoregressive planning (i.e., commit to the first sample from our policy module) instead of our proposed simulation-based planning with world model. Formally, the planning process is simplified to the following:
    $a_{t}^{\prime*}=\operatorname*{arg\,max}_{a_{t}^{\prime}}p_{\tilde{\pi}}(a_{t}^{\prime}\mid\hat{s}_{t})$
Figure 6: Overview of performance comparison between SimuRA and baselines. The full architecture shows clear advantage over the baseline BrowsingAgent, improving the performance on complex website navigation from 0% to 32.2%. Our proposed world model reasoning for planning also consistently improves over simple planning with autoregressive LLM by up to 124%. *We implement the autoregressive planner with o3-mini and test on the complex website navigation dataset.
Implementation for Web Browsing
Figure 5 presents our implementation when applied to web browsing. We use prompts tailored to the web environments in this example, but plan to extend to other environments and move towards training a single agent model that can act optimally in a wider range of environments and tasks, which is an exciting next step. At each step $t$, the agent receives the observation $o_{t}$ as the HTML-based accessibility tree visible through the browser’s viewport (an example is provided in Appendix A). The agent then uses encoder LLM $h$ to summarizes the observation as $\tilde{s}_{t}\sim p_{h}(\cdot\mid o_{t})$, and then add it to a selective memory of past summaries and simulated actions $\{m(\tilde{s}_{k},a_{k}^{\prime*})\}_{k=1}^{t-1}$ to form the estimated world state $\hat{s}_{t}=[m(\tilde{s}_{1},a_{1}^{\prime*}),\dots,m(\tilde{s}_{t-1},a_{t-1}^{\prime*}),\tilde{s}_{t}]$ for planning. During planning, we sample $M$ simulated actions $a^{\prime}_{t}$ from the policy $\tilde{\pi}$, cluster them into distinct actions, and use the world model $f$ to predict the next summary as $\tilde{s}_{t+1}\sim p_{f}(\cdot\mid\hat{s}_{t},a_{t}^{\prime})$ to form the next state $\hat{s}_{t+1}=[m(\tilde{s}_{1},a_{1}^{\prime*}),\dots,m(\tilde{s}_{t},a_{t}^{\prime}),\tilde{s}_{t+1}]$; this repeats until the planning horizon $T$. To evaluate the terminal state $\hat{s}_{T}$ with critic $v$, we prompt the LLM to generate categorical answers and convert them into numerical scores (e.g., “success” receives a score of 1), and repeat for $N$ times to capture the fine-grained differences between states. Following previous work koh2024tree ; gu2025llmsecretlyworldmodel , we set $M=N=20$ and $T=t+1$, and use DFS as the search algorithm. We implement the planning process using LLM Reasoners hao2024llm , a library for LLM-based complex reasoning using advanced algorithms. After the planner selects the simulated action $a_{t}^{\prime*}$, we update the memory with $m(\tilde{s}_{t},a_{t}^{\prime*})$. For the actor $\alpha$, we additionally include the observation text $o_{t}$ in the prompt to ensure the action grounding. All the prompts are included in Appendix B.
Overview of Results
An overview of our results is presented in Figure 6. Across all 3 categories of tasks, our architecture shows a clear advantage over the baseline BrowsingAgent, specifically increase the success rate on complex website navigation from 0% with a popular open-web agent baseline to 32.2%. Our proposed world model reasoning for planning also consistently improves over a matched autoregressive reasoning baseline by up to 124% in task-completion rate. Perhaps surprisingly, autoregressive planning with reasoning LLMs like o3-mini have close-to-zero success rate in Complex Website Navigation, showing that even a strong textual-reasoning model can have limited success with planning for web interactions. All experiments were conducted between November and December 2024 using the best publicly available models and browser environments at that time, except for the o3-mini experiment which was performed in February 2025. While subsequent model updates may yield higher absolute scores on these tasks, our results demonstrate that explicit world-model simulation provides a measurable and repeatable advantage over purely autoregressive planning under identical conditions. In the subsections below, we will introduce the evaluation settings and discuss the results in more detail.
4.1 Complex Website Navigation
A distinguishing feature of web agents is the ability to gather live information (e.g., flights, stock prices, social media) not present in the training data of foundation models due to its rapid update (yoran2024assistantbench, ). For many questions (e.g., the price of the earliest flight tomorrow), LLMs without such grounding often result in hallucinations (see Figure 7 for an example). In practice, however, obtaining the information is challenging, as many websites are very complex and difficult to navigate (e.g., execute a flight search query on a travel website and filter through the results), which calls for superb reasoning skills on the part of the agent.
Dataset
Existing benchmarks he2024webvoyagerbuildingendtoendweb have contributed valuable infrastructure for evaluating web agents in diverse real-world settings. However, these datasets are primarily constructed through self-instruct–based task generation followed by human verification, which, while ensuring breadth and realism, offers limited control over task structure and constraint complexity. Therefore, the resulting tasks may follow distributions favored by the language model used for generation, making it difficult to systematically assess how an agent’s reasoning performance scales with task difficulty or compositional complexity. To address this limitation, we introduce FlightQA, a dataset designed for controlled and scalable evaluation of reasoning robustness in live web environments. We focus on flight search, a representative real-world use case that requires multi-step reasoning and precise constraint handling. In this setup, the user requests a flight satisfying a list of constraints (e.g., one-way, from New York to Los Angeles), and the agent must operate an online flight search interface to locate and return a valid option. To systematically evaluate the agent’s reasoning ability, we construct questions with varying number of constraints by iteratively adding to the list, which enables a type of counterfactual analysis that controls for the confounding effect of specific constraint configurations. For instance, an agent capable of robust reasoning should succeed when an additional constraint is added to an existing query, whereas one relying on rote memorization will likely fail.
We illustrate our data collection process in Figure 8. To ensure scalability and controllability, we prompt a LLM to first generate a list of $C$ starting constraints, repeating for $N$ times. After that, we prompt the LLM to iteratively add constraints to the lists one at a time, repeating for $K$ times. Finally, we prompt the LLM to convert each constraint list into a question in natural language. In practice, we set $C=3$, $N=15$, and $K=5$, which results in FlightQA, a dataset consisting of 90 questions with 15 sequences of constraint lists where the number of constraints increases from 3 to 8. We use gpt-4o to perform all the data generation steps. The initial question generation and question expansion prompts are included in Appendix C
Figure 7: Faced with the question “What is the earliest-arriving flight tomorrow from Pittsburgh to Zurich?” ChatGPT-4o browsed the frontpage of Kayak.com and hallucinated a flight that arrives at 10:45am on the following day as the answer (left). Performing the search on Kayak.com, however, shows that the earliest-arriving flight lands in Zurich at 6:10am on the next day (right). The question was asked on December 17th, 2024.
Evaluation
Because FlightQA involves querying live information from the open internet, it is impossible to establish ground truth answers due to the constantly evolving flight pricing and availability. Inspired by previous work on evaluation for generated text (deng2021compression, ), we propose to evaluate the agent response based on two quality aspects: groundedness for whether the response is supported by the interaction history and relevance for whether the response satisfies user constraints to the extent allowed by the results (e.g., if the search results do not include any flight that satisfies all user constraints). Due to the strong ability of LLMs in evaluating generated text (liu-etal-2023-g, ), we prompt LLMs to assess the two quality aspects of the agent response. Specifically, we include all browser observations in the agent’s trajectory over $T$ steps $(o_{1},o_{2},…o_{T})$, the constraint list, the question, and the agent response, and ask the LLM to provide judgment on the groundedness and relevance of the response. We further define an answer to be correct when it is both grounded and relevant. We also include the evaluation prompt in Appendix C.
Experiment Setup
We ran the experiments and evaluation using gpt-4o between November 24th, 2024 and December 9th, 2024. To evaluate the capabilities of autoregressive reasoning LLMs trained with reinforcement learning, we also included variants of the autoregressive reasoning baseline where the planner is implemented using o1 openai_learning_to_reason and o3-mini openai_o3_mini_2025 and ran them between February 3rd to 5th, 2025. For the environment, we use BrowserGym (workarena2024, ), a popular open-source browser sandbox. We stop each run when the agent provides a response or after the agent takes 30 actions, whichever comes first. We also mark the run as failed when the agent repeats the same action for 3 times consecutively or when the agent causes more than 3 errors while interacting with the browser.
Figure 8: Illustration of the data generation process for the FlightQA dataset. We first prompt a LLM to generate NN lists of $C$ starting constraints (Constraint Generation). Then, we prompt the LLM to iteratively add constraints to the lists one by one, repeating for $K$ times (Constraint Extension). Finally, we prompt the LLM to convert each constraint list into a question in natural language (Question Generation).
Results
We present our Complex Website Navigation results in Table 1. Compared to BrowsingAgent which fails completely in this task, our full architecture improves the accuracy from 0% to 32.2%. Within our architecture, our proposed world-model-based planning shows superior performance over autoregressive reasoning with a 124% improvement (significant at the 0.01 level). The other components in our architecture, which communicate using the structured latent space of model-generated natural language (e.g., observation summary and selective memory), also result in more coherent behavior by reducing the action error rate in BrowsingAgent from 93.3% to 1.1%. However, the autoregressive reasoning still results in frequent repetitions, which is mitigated by world model-based planning (44.4% $\rightarrow$ 18.9%). Perhaps surprisingly, o1 and o3-mini receive close to 0% success rates when serving as autoregressive planners, suggesting that even strong textual-reasoning models can have limited success with planning for web interactions. We do not include BrowsingAgent with o1 and o3-mini as the resulting agent frequently hallucinate answers without interacting with the webpage, which precludes them as viable agents. Those hallucinations, however, fool the LLM evaluator at significant rates, suggesting additional work is needed to secure its robustness.
Performance (%) Outcomes (%)
Method Correct Grounded Relevant
Response
Returned
Browser
Crashed
Max Steps
Reached
Repetitive
Actions
Action
Errors
OpenHands BrowsingAgent 0.0 0.0 0.0 0.0 3.3 3.3 0.0 93.3
SimuRA (Ours)
Autoregressive Planning 14.4 15.6 14.4 16.7 0.0 37.8 44.4 1.1
– with o1 1.1 1.1 1.1 1.1 11.1 40.0 37.8 10.0
– with o3-mini 3.3 4.4 3.3 4.4 3.3 51.1 32.2 8.9
World Model Planning 32.2** 36.7 32.2 38.9 1.1 40.0 18.9 1.1
Table 1: Performance and outcome statistics for the FlightQA dataset. Our architecture increases the accuracy from 0% in OpenHands BrowsingAgent to 32.2%. Reasoning by world model simulation also clearly outperforms autoregressive reasoning by 124%. ** indicates being significantly higher than the second-best method at the statistical significance level of 0.01 ($p<0.01$) based on pairwise t-test. We implement the autoregressive planner with o1 and o3-mini, respectively.
Analysis of Reasoning Ability
To compare the reasoning abilities of autoregressive and world-model planners within our architecture, we visualize the percentage of correct responses vs number of constraints in Figure 9. As the questions in FlightQA are generated based on iteratively expanded constraint lists, this analysis should faithfully reflect the effect of increasing constraints while controlling for other confounders such as specific constraint sets. Based on our data samples, world model planning shows consistent advantage over autoregressive planning as we increase the number of constraints, showing signs of improved reasoning ability. The performance for both methods decreases with more constraints initially but then increases sharply at 7 constraints before dropping again, which may reflect memorization in the backend LLM or implicit constraints in questions with fewer explicit constraints.
Figure 9: % correct and % response returned vs. number of constraints for FlightQA. Based on our data samples, world model planning consistently outperforms autoregressive planning as we increase the number of constraints, showing signs of improved reasoning ability.
4.2 Multi-Hop, Multi-Website QA
Another type of challenging questions for web agents is those that require gathering information about multiple entities over multiple websites. For instance, given the question “What are the availabilities of the top-10 restaurants in Paris for a dinner next week?”, an agent must first find the top-10 restaurants in Paris, then look up the availability of each restaurant, and finally compile the information into a response to the user. Whereas complex website navigation stresses the depth of individual websites, multi-hop, multi-website QA concerns the breadth of websites to navigate over long-horizon interactions.
Dataset
To evaluate agent abilities for multi-hop, multi-website QA, we adopt the FanOutQA (zhu-etal-2024-fanoutqa, ) dataset, which consists of questions of exactly this nature. Due to resource constraints, we evaluate on the first 100 examples of the dev set. As the results show, however, the smaller sample size is sufficient to show statistically significant differences between methods.
Experiment Setup
We ran the experiments using gpt-4o-2024-05-13 between November 10th, 2024 and December 8th, 2024. We noticed that our architecture with world-model-based planning deteriorates in performance when using the newer versions of gpt-4o, which may be due to additional training which changed the model’s response patterns to the same prompts. We operate the browser using the same rules as in experiments for Complex Website Navigation.
Results
We present our results on Multi-Hop, Multi-Website QA in Table 2. Again, our method increases the accuracy from 17.0% to 29.8% and world model planning improves over autoregressive planning by 48.6% (p-value = 0.011). BrowsingAgent achieves fair performance even though it cannot memorize information from different websites, often due to some questions in the dataset being answerable based on information from a single Wikipedia page (e.g., What are the publication dates for all of the Harry Potter books?). Despite this, our architecture improves over BrowsingAgent even without world model planning by dramatically reducing action errors (43% $\rightarrow$ 10%). Browser crashes make a sizable contribution to agent failures (24% for our architecture), indicating room for improvement in the tooling for open-web navigation.
Performance (%) Outcomes (%)
Method Acc. Acc. (Strict)
Response
Returned
Browser
Crashed
Max Steps
Reached
Repetitive
Actions
Action
Error
Parsing
Error
OpenHands BrowsingAgent 17.0 4.0 32.0 17.0 8.0 0.0 43.0 0.0
SimuRA (Ours)
Autoregressive Planning 20.2 3.0 37.0 24.0 10.0 18.0 10.0 1.0
World Model Planning 29.8* 4.0 55.0 24.0 12.0 8.0 1.0 0.0
Table 2: Performance and outcome statistics for the FanOutQA dataset. Acc. (Strict) refers to the percentage of responses that exactly match the groundtruth. Our architecture clearly outperforms the baseline BrowsingAgent. Reasoning by world model increases the response rate and fact-level accuracy vs. autoregressive planning by 48.6% and 47.5%, respectively. * indicates being significantly higher than the second-best method at the 0.05 level based on pairwise t-test.
4.3 General Web Automation
Last but not least, web agents are often tasked with performing various work tedious to human users (e.g., online shopping, managing social media). These tasks often require the ability to interact with a range of websites of moderate complexity. As an example, given the question “Summarize customer reviews for Amazon Echo Dot 3rd generation,” the agent should navigate a shopping website to locate and go over all the customer reviews of said product before summarizing the content for the user.
Dataset
To evaluate general web automation capabilities, we adopt the WebArena (zhou2023webarena, ) benchmark, a standard environment for testing web agents which features a range of simulated websites including a Reddit-like social forum, a shopping site, a GitLab-based code management platform, a map, and a Wikipedia-like encyclopedia. Following the evaluation for Multi-Hop, Multi-Website QA, we take a random subset of 100 examples.
Experiment Setup
We run the experiments using gpt-4o over BrowserGym accessed via the OpenHands platform which provides a uniform evaluation procedure. Because WebArena demands a specific response format for evaluation, we rewrote the agent description to steer the agent answer format accordingly (Appendix B.1). We keep all other environment rules the same as previous experiments, except for setting the maximum allowed steps to 15 which is consistent with the default setting of WebArena.
Results
We present our results on General Web Automation in Table 3. Continuing the patterns from previous experiments, our proposed architecture improves over BrowsingAgent by up to 91.7%, while within our architecture, world model reasoning improves over autoregressive reasoning by 21.1%, highlighting the comparative advantage under the given experimental setup. Due to following the environment and evaluator provided by OpenHands, which prioritizes open-web browsing and differs from the standard benchmark setup, the success rates are not directly comparable to those reported in prior work zhou2023webarena . Nevertheless, these results demonstrate the relative advantage of our proposed simulative reasoning architecture.
Method Success Rate (%)
OpenHands BrowsingAgent 12.0
SimuRA (Ours)
Autoregressive Planning 19.0
Ours (World Model Planning) 23.0
Table 3: Results on a random 100-sample subset of WebArena. Our architecture improves over BrowsingAgent by up to 91.7%, while world model planning improves over autoregressive planning by 21.1%.
5 Limitations
Due to the modular pipeline and thorough exploration of multiple plans in world-model-based simulation, the current agent takes longer than typical reactive agents to run. Speeding up world-model-based reasoning with appropriate caching and parallelization strategies is an important part of our future work. In addition, agent capabilities can be limited by the tooling. For example, with open-source browser environments, web agents are often blocked by Captcha or anti-scraping tools from certain websites. Deeper integration with user browser can help solve this issue. As agent-based automation become more integrated into browsing and computer-use workflows, we also encourage conversations around fair use and protocols around agent access of certain websites. We are currently only using the text portion of the webpage observations, which can miss information like images and layout information (e.g., occlusions). While existing work are experimenting with visual-based web browsing, it is still challenging to combine multimodal perception and planning, which we are excited to keep working on.
6 Conclusion
In this paper, we have presented SimuRA, a goal-oriented architecture for agentic reasoning and decision-making. By augmenting autoregressive reasoning with simulation-based planning through a world model and by representing internal belief states in the semantically structured latent space of natural language, SimuRA demonstrates consistent improvements across a range of web-interaction tasks. In particular, our results suggest that explicit world-model reasoning can enhance the planning and reasoning capacity of agents beyond what purely autoregressive reasoning provides, underscoring the value of simulative reasoning as a core component of general intelligence.
Looking ahead, we see SimuRA as a foundation for developing more capable, robust, and controllable agentic systems. On the capability side, we aim to test on more types of environments (e.g., embodied sandboxes and physical space) and to continue developing functional components that including multi-agent interaction and long-term memory. On the safety and alignment side, we look forward to engaging the community in discussions about how explicit world-model reasoning may help build agents that remain transparent and aligned with human values and collective welfare.
Acknowledgment
This work was supported in part by the Samsung GRO Project “Efficient Designs for Generative and Agent LLM Development.” We thank Graham Neubig and Zora Wang from NeuLab; Yilin Shen and Hongxia Jin from Samsung Research; Zhoujun Cheng, Shibo Hao, and Xinyu Pi from MixLab; Han Guo, Nicholas Ho, and Bowen Tan from SAILING Lab; Li Erran Li from AWS, and Sarah Cheah and Hector Ren from MBZUAI for their insightful feedback and discussions. We are also grateful for their helpful suggestions throughout the project. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Samsung.
Appendix A Details on Web Browsing Environment
Example Observation
URL https://www.google.com/travel/flights
Scroll Position: 0, Window Height: 720, Webpage Height: 3024, Remaining Pixels: 2304, Scrolling
Progress: 23.8%
RootWebArea ’Google Flights - Find Cheap Flight Options & Track Prices’
    [149] banner ’’
        [160] button ’Main menu’, clickable, expanded=False
            [161] image ’’
        [168] link ’Google’, clickable
        StaticText ’Skip to main content’
        StaticText ’Accessibility feedback’
        [186] navigation ’’
            [189] link ’Travel’
                [193] image ’’
            [197] link ’Explore’
                [201] image ’’
            [205] link ’Flights’
                [209] image ’’
            [213] link ’Hotels’
                [217] image ’’
            [221] link ’Vacation rentals’
                [225] image ’’
        [235] button ’Change appearance’, hasPopup=’menu’, expanded=False
            [240] image ’’
        [249] button ’Google apps’, clickable, expanded=False
            [250] image ’’
        [251] link ’Sign in’, clickable
    [342] image ’’
    StaticText ’Flights’
    [346] search ’Flight’
[…]
              text’, hasPopup=’listbox’, expanded=False, controls=’i9’
            [364] image ’’
        [399] button ’1 passenger, change number of passengers.’, hasPopup=’dialog’
            [404] image ’’
            [406] image ’’
[…]
              text’, hasPopup=’listbox’, expanded=False, controls=’i22’
                [529] image ’’
        [576] combobox ’Where from?’ value=’Pittsburgh’, clickable, autocomplete=’inline’,
              hasPopup=’menu’, expanded=False
        [580] image ’’
        [628] button ’Swap origin and destination.’, disabled=True
                [631] image ’’
        [638] combobox ’Where to?’, clickable, focused, autocomplete=’inline’, hasPopup=’menu’,
              expanded=False
        [641] image ’’
        generic ’’, hidden=True
        [690] image ’’
        [691] textbox ’Departure’, clickable, describedby=’i32’
        [712] textbox ’Return’, clickable, describedby=’i32’
        generic ’’, hidden=True
        [857] button ’Explore destinations’
    [866] Section ’’
        [867] heading ’Find cheap flights from Pittsburgh to anywhereMore information on suggested
              flights.’
            [871] button ’More information on suggested flights.’, hasPopup=’menu’
                [873] image ’’
        [904] list ’’, clickable
            [905] listitem ’’
                StaticText ’Pittsburgh’
            [907] listitem ’’
                [908] button ’Cleveland’
            [909] listitem ’’
                [910] button ’Columbus’
            [911] listitem ’’
                [912] button ’Akron’
    StaticText ’San Francisco’
    StaticText ’\\$128’
    StaticText ’Jan 9 — Jan 16’
    StaticText ’1 stop’
    StaticText ’·’
    StaticText ’10 hr 30 min’
    StaticText ’New York’
    StaticText ’\\$68’
    StaticText ’Dec 7 — Dec 14’
Appendix B Prompts for Web Browsing Implementation
Prompt for Agent Identity # Name:
Web Browsing Agent
# Description:
An information and automation assistant who responds to user instructions by browsing the internet. The assistant strives to answer each question accurately, thoroughly, efficiently, and politely, and to be forthright when it is impossible to answer the question or carry out the instruction. The assistant will end the task once it sends a message to the user.
# Observation Space:
The text representation and screenshot of the part of webpage visible in the viewport of a browser. Here is an abstract description of the information available in the webpage text representation:
- Identification Information:
- URL: The web address that specifies the location of the webpage. - Document Properties: Attributes such as scroll position and viewport dimensions that describe the current viewing context. - Structural Hierarchy: - Root Element: The primary container for the webpage, indicating its overall theme or purpose. - Nested Elements: A hierarchy of sections, containers, and components that organize content logically (e.g., headers, footers, sidebars). - Interactive Components: - Links: Elements that can be clicked to navigate to other pages or sections, often labeled descriptively. - Buttons: Interactive controls that trigger actions (e.g., submitting forms, opening menus). - Content Types: - Text: Main content, headings, and subheadings that provide information and context. - Images and Media: Visual elements that enhance the understanding or appeal of the content. - Forms and Inputs: Fields for user input, including text boxes, dropdowns, and checkboxes. - Functional Areas: - Navigation Menus: Organized sets of links that allow users to explore different sections of the site. - Search Interface: Components that enable users to search for content within the site, including input fields and associated buttons. - State Information: - Visibility and Expand/Collapse States: Indicators showing whether certain elements are active, visible, or in a collapsed state, impacting user interaction. - Focus States: Information on which elements are currently focused, important for keyboard navigation and accessibility.
Prompt for Agent Identity (Continued) - Accessibility Features: - Role and Description Information: Metadata that provides context about the purpose of elements, useful for screen readers and assistive technologies. - General User Considerations: - Navigation: Recognizing how to traverse the webpage using links and buttons. - Interactivity: Understanding how to engage with forms, search fields, and dynamic components. - Content Engagement: Identifying and interpreting various content types to glean necessary information. # Action Space:
13 different types of actions are available.
noop(wait_ms: float = 1000)
[…]
Prompt for Agent Identity (Continued) hover(bid: str) Examples: hover(’b8’) press(bid: str, key_comb: str) Examples: press(’88’, ’Backspace’) press(’a26’, ’Control+a’) press(’a61’, ’Meta+Shift+t’) focus(bid: str) Examples: focus(’b455’) clear(bid: str) Examples: clear(’996’) drag_and_drop(from_bid: str, to_bid: str) Examples: drag_and_drop(’56’, ’498’) upload_file(bid: str, file: str | list[str]) Examples: upload_file(’572’, ’my_receipt.pdf’) upload_file(’63’, [’/home/bob/Documents/image.jpg’, ’/home/bob/Documents/file.zip’]) Only a single action can be provided at once. Example: fill(’a12’, ’example with "quotes"’)
# Instruction: {user_instruction}
# Current Date and Time: {current_datetime}
Prompt for Encoder # Observation:
{observation}
# State:
Describe all the elements in the current webpage observation. Note any dialog, progress indicators, or error messages. Include any interactive elements and their values or if they are blank. Note any detailed information such as facts, entities, or data that are relevant to the task. Report any error messages like whether the last action was correct. Try to be as comprehensive and detailed as possible.
Wrap your response in the tag <state> and </state>.
Prompt for Policy {memory}
# Current State:
{state}
# Intent:
Describe the action the assistant should take next to carry out the user’s instruction. Avoid using phrases such as "To accomplish the goal," "I will," "To proceed.". Avoid ending with phrases like "to execute the search." Describe one action at a time and avoid combining multiple steps. Refrain from mentioning specific element IDs as they may change during execution. Limit your response to one phrase and include any details that help select the correct action. Be creative and propose novel methods to achieve the goal. Avoid creating accounts without user permission or providing personal information. Concrete example would be "Go to the home page of Google Flights." and "Click on the ’Search’ button."
Wrap your response in the following format:
<think>
Your thoughts and reasoning process
</think>
<intent>
Description of the action to perform next
</intent>
Prompt for World Model {memory}
# Current State:
{state}
# Memory Update:
{memory_update}
# Action Intent:
{plan}
# Next State:
Describe all the elements in the webpage after the agent attempts to carry out the intent. Note that the execution may not be successful, so you will have to infer the result of the action. Note any dialog, progress indicators, or error messages. Include any interactive elements and their values or if they are blank. Note any detailed information such as facts, entities, or data that are relevant to the task. Report any error messages displayed. Try to be as comprehensive and detailed as possible.
Wrap your response in the following format:
<next_state>
Follow the format of the current state description. Use present tense. Avoid starting phrases like "Based on the interaction history, current state, and current intent".
</next_state>
Prompt for Critic {memory}
# Final State:
{state}
# Task Success and Progress:
Your task is to evaluate the performance of the agent. Given the agent’s instruction, interaction history, the final state of the webpage, and the agent’s responses to the user if any, your goal is to decide whether the agent’s execution is successful or not. If the current state is a failure but it looks like the agent is on the right track towards success, you should also output as such.
Wrap your response in the following format:
<think>
Your thoughts and reasoning process
</think>
<status>
"success" or "failure"
</status>
<on_the_right_track>
"yes" or "no"
</on_the_right_track>
Prompt for Memory Update {memory}
# State:
{state}
# Action Intent:
{plan}
# Memory Update:
Summarize the changes in the webpage observation that should be remembered for achieving your goal and for predicting the next state. Note any new elements, any elements no longer visible, or any changes in the content of existing elements. Also note if there is no change. Include any other inferred information that may help you decide the next action, such as whether an action intent is successful, or whether progress has been made or reversed. Do not include your next planned actions. Revise your belief from previous history if the current state contradicts it.
Wrap your response in the tag <memory_update> and </memory_update>.
Prompt for Actor {memory}
# Observation:
{observation}
# Current State:
{state}
# Current Intent:
{plan}
# Action:
Choose an API call that will carry out the intent when executed in the webpage. Use only one action at a time. You must not enclose bid inputs in [brackets] but instead in ’single quotes’. Interact only with elements in the current step observation. Your response will be executed as a Python function call, so ensure it adheres to the format and argument data type specifications defined in the action space.
Wrap your response in the tag <action> and </action>.
Prompt for Action Clustering Here is the action space for a browser agent to navigate in a webpage:
16 different types of actions are available:
noop(wait_ms: float = 1000)
send_msg_to_user(text: str)
scroll(delta_x: float, delta_y: float)
fill(bid: str, value: str)
select_option(bid: str, options: str | list[str])
click(bid: str, button: Literal[’left’, ’middle’, ’right’] = ’left’, modifiers: list[typing.Literal[’Alt’, ’Control’, ’Meta’, ’Shift’]] = [])
dblclick(bid: str, button: Literal[’left’, ’middle’, ’right’] = ’left’, modifiers: list[typing.Literal[’Alt’, ’Control’, ’Meta’, ’Shift’]] = [])
hover(bid: str)
press(bid: str, key_comb: str)
focus(bid: str)
clear(bid: str)
drag_and_drop(from_bid: str, to_bid: str)
upload_file(bid: str, file: str | list[str])
go_back()
go_forward()
goto(url: str)
Only a single action can be provided at once. Example: fill(’a12’, ’example with "quotes"’)
Below, you will find lists of intents, or natural language descriptions of actions that, when executed, will translate to one of the function calls above. The intents will be provided in the following JSON format:
‘‘‘json
  "intent_id": "intent description"
Your task is to cluster list of intents into semantically equivalent groups, where each group represents intents that lead to the same action when executed (i.e., navigating to the Google homepage is translated to goto(’https://www.google.com’)) and would therefore correspond to the same API call in a Playwright browser. Intents that use different wording but convey the same action should be grouped together. Try to minimize the number of clusters.
Prompt for Action Clustering (Continued) Represent the clustering results using a JSON object where each cluster has a unique identifier, and each identifier maps to a list of actions in that cluster. See below for an abstract example:
‘‘‘json
  "cluster_id": {
    "intent": "representative intent name for this cluster",
    "candidates": [
      "<list of intent ids that belong to this cluster>
Concrete Example 1:
Dictionary of Intents:
‘‘‘json
  "0": "Navigate to the Google homepage by entering its URL.",
  "1": "Go to the Google homepage.",
  "2": "Go to the Google homepage",
  "3": "Go to the Google homepage by navigating to
        ’https://www.google.com’",
  "4": "Go to the home page of Google"
["Navigate to the Google homepage by entering its URL.", "Go to the Google homepage.", "Go to the Google homepage", "Go to the Google homepage by navigating to ḧttps://www.google.com"̈, "Go to the home page of Google"]
Clustering Results:
‘‘‘json
  "cluster_1": {
    "intent": "Navigate to the Google homepage",
    "candidates": [0, 1, 2, 3, 4]
Concrete Example 2:
Dictionary of Intents:
{action_candidate_json}
Clustering Results:
B.1 Adaptation for WebArena Evaluation
Agent Description for WebArena Evaluation An information and automation assistant that interacts with the browser and responds to user instructions. The response follows the following rules: 1. When the intent is a question, and a complete answer to the question has been found, then send the answer to the user; 2. the intent wants to locate specific information or navigate to a particular section of a site, and the current page satisfies, then stop and tell the user you found the required information; 3. the intent want to conduct an operation, and has been done, then stop and tell the user the operation has been completed. The assistant should try to achieve the goal in the current site without navigating to sites like Google. Be forthright when it is impossible to answer the question or carry out the task. The assistant will end the task once it sends a message to the user.
Appendix C Prompts for Generating and Evaluating on the FlightQA Dataset
Prompt for Generating Initial Constraints and Questions System: You are a creative writer who is an expert at crafting questions to help train assistants who answer user queries. Current date and time: {current_datetime} Instruction: Your task is to create a robust benchmark for evaluating an AI’s ability to search for flights through a platform like Google Flights. To ensure the dataset effectively represents real-world use cases. Here are some important factors to consider: 1. Diversity of Queries - Range of Destinations: Include both common and obscure destinations to test how well the model handles varying levels of demand. - Dates and Duration: Include different date ranges, including last-minute flights, peak travel dates (like holidays), and extended trips. Ensure there’s a variety in trip duration as well. - Passenger Variability: Include solo travelers, families, and group travel (e.g., one adult vs. two adults and two children) since these change the search parameters and price results. - Class and Preference: Vary preferences like cabin class (economy, business, first) and filters (non-stop, one stop, preferred airlines, etc.). - Budget Constraints: Test price sensitivity by setting different budget limits to see how well the AI handles trade-offs. 2. Complexity of Requirements - Multi-Leg Flights: Add queries for multi-city trips or those requiring complex layovers. - Dynamic Constraints: Include queries with dynamic constraints, such as “find the cheapest flight but depart between 8-10 AM,” to see if the model can adapt its search within specific time frames. - Conditional Preferences: Test cases where users might want options based on multiple conditions, like “either find the cheapest non-stop or the shortest two-stop option.” In practice, the questions typically vary in the following dimensions: - Ticket type (round-trip, one-way, etc.) - Routes (origin and destination) - Layover location(s) - Dates (departure and/or return) - Flight time (departure and arrival) - Total flight time - Airlines - Cabin class (economy, business, etc.) - Aircraft - Eco-friendly options (CO2 Emissions) Given a number of constraints, you should first provide a list of constraints, with the number of constraints equal to the specification. After that, you will generate a question a typical user will ask which imposes those constraints. You should repeat this for at least 7 times to generate a set of questions with simple language. Make sure that the number of constraints in the question matches the number of constraints specified. Do not include constraints about the number of passengers. If the constraint is a date, you can use relative dates (e.g., "tomorrow", "next month", "after 8 PM", etc.). Avoid using specific dates like "December 25th" to ensure the questions are relevant throughout the year. Your response should follow the JSON format below:
Prompt for Generating Initial Constraints and Questions (Continued) Number of Constraints: <num_constraints>
    "num_constraints": <num_constraints>,
    "questions": [
            "constraints": [<constraints>],
            "question": <question>
Below is a concrete example: Number of Constraints: 3
    "num_constraints": 3,
    "questions": [
            "constraints": ["one-way", "New York to London",
                            "departing next Friday"],
            "question": "Can you find a one-way flight from New York
                         to London departing next Friday?"
Prompt for Iteratively Expanding Constraints and Questions System: [Same as above] Instruction: [Same as above until “Your response should follow”] Your response should follow the JSON format below: Maximum number of constraints: <max_constraints> Starting constraints and questions:
    "num_constraints": <num_constraints>,
    "constraints": [<constraints>],
    "question": <question>
Questions with increasing complexity:
    "questions": [
            "num_constraints": <starting num_constraints + 1>,
            "constraints": [<previous constraints with 1 additional>],
            "question": <question>
            "num_constraints": <starting num_constraints + 2>,
            "constraints": [<previous constraints with 2 additional>],
            "question": <question>
        ... (continue until reaching the maximum number of constraints)
Your Response: Maximum number of constraints: {max_num_constraints} Starting constraints and questions: {starting_constraint_questions} Questions with increasing complexity:
Prompt for Evaluation # Interaction Date and Time: {interaction_datetime} # Interaction History:
[Concatenation of observations from all steps]
Above are the webpages an assistant interacted with while trying to answer the user’s query. The user is looking for flights with the following constraints: {constraints} Here is the exact query provided by the user: {goal} Here is the assistant’s response: {message} Your task is to evaluate two aspects of the response: 1) Whether the assistant’s response is supported by the interaction history, and 2) Whether the assistant’s response satisfies the user constraints to the extent allowed by the results. Some Context: - To determine the seating class of a flight being returned, refer to the value of the "Change seating class" combobox. - It is not always possible to satisfy all the user constraints. In this case, examine whether the response is as close to the user constraints as possible. Answer in the following format: <think> Your thoughts and reasoning. </think> <grounding> Your assessment of whether the response is supported by the interaction history. Answer "yes" or "no" </grounding> <relevance> Your assessment of whether the response satisfies the user constraints to the extent allowed by the results. Answer "yes" or "no" </relevance>